OpenStudy (anonymous):

Fill in the blank for the explicit rule that produces this series. -6, -3, 2, 9, 18, ... a[n] = n^2 – _____

myininaya (myininaya):

since the second common differences are a constant then your sequence should be in the form a[n]=An^2+Bn+C it looks like they already give you A=1 so a[n]=n^2+Bn+C we can find B and C be using the numbers given in your sequence use (1,-6) and (2,-3) two give you two equations these 2 equations will allow you to solve for B and C

OpenStudy (anonymous):

Thank you (:

myininaya (myininaya):

np

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{-6, -3, 2, 9, 18, ........\\~\\ 1-7, 4-7-, 9-7, 16-7, 25-7, ........\\~\\ 1^2-7, 2^2-7-, 3^2-7, 4^2-7, 5^2-7, ........\\~\\ }\end{align}\)